OpenStudy (anonymous):

Paul needs 2 1/4 yards of fabric to make a tablecloth. How many tablecloths can he make from 18 1/2yards of fabric? A. 72 tablecloths B. 10 tablecloths C. 9 tablecloths D. 8 tablecloths

OpenStudy (anonymous):

2 1/4 = 9/4 ; 18 1/2 = 37/2 = 74/4 1 table cloth = 9/4 yards Total yards = 74/4 Divide and you get 8 tablecloths with 1/2 yards remaining.